Kaihalojr:

Simplify. (4x - 12)/(x^2 - 9) * (x+3)/(5x^2-25x) / (28)/(x^2-25)

Kaihalojr:

I need to factor out like terms

Kaihalojr:

also These are fractions.

Hero:

Okay, so @Kaihalojr, what does \(4x - 12\) factor to?

Kaihalojr:

@Hero it can factor to x - 3 right?

Hero:

Yes, but what did you factor out in order to get \(x - 3\) as one of the factors?

Kaihalojr:

4 because that's a number that they are both divisible by.

Hero:

Okay, so wouldn't that make both \(4\) and \(x - 3\) factors of \(4(x - 3)\)?

Hero:

Oops, wrote that wrong. Guess that's a freebee for you.

Hero:

So yeah \(4x - 12\) factors to \(4(x - 3)\) since \(4(x - 3) = 4x - 12\)

Kaihalojr:

so for the bottom half of the fraction isn't x² - 9 a difference of squares?

Hero:

Correct and how would you write that in factored form?

Kaihalojr:

x - 3

Kaihalojr:

so this whole fraction is basically x - 3/x - 3

Hero:

Is it?

Kaihalojr:

wait no wouldn't it actually be 4(x-3)/(x-3)²?

Hero:

Yes, but doesn't \(\dfrac{x - 3}{x - 3} = 1\) And knowing that couldn't we just re-write \((4(x - 3))/(x - 3)^2\) as \(\dfrac{4}{x - 3} \cdot\ \dfrac{x - 3}{x - 3}\) and then as \(\dfrac{4}{x - 3} \cdot\ 1\)

Kaihalojr:

I think so yes

Hero:

Okay, so that should tell you what the first fraction should be

Hero:

Any ideas on how to factor the second fraction?

Kaihalojr:

I think the upper half doesn't need to be factored right? just the bottom?

Hero:

Are you sure about that?

Kaihalojr:

no that's why I'm asking you

Hero:

Okay, so basically the denominators need to be factored out in this case.

Hero:

Let's write it out

Kaihalojr:

I'm thinking we divide 25 and 5 by 5 so it's 1 and 5 then we divide x^2 and x by x so it's x and 1 correct?

Hero:

\(\dfrac{\dfrac{x + 3}{5x^2 -25x}}{\dfrac{28}{x^2 - 25}}\)

Hero:

What can we factor from the denominator of the top fraction?

Kaihalojr:

oh we can divide the whole thing by x right so it becomes 3/ 5x - 25?

Hero:

What can we FACTOR from the denominator of the top fraction?

Kaihalojr:

5 right because 5 and 25 can each be factored by 5 or x for x^2 and x

Hero:

You can factor \(5x\) not \(5\) or \(x\).

Kaihalojr:

oh yeah that makes sense so then it's x - 5 right?

Hero:

Yes, and you know the denominator of the bottom fraction is difference of squares.

Hero:

By the way, the first fraction is \(\dfrac{4}{x + 3}\)

Kaihalojr:

which multiplied by (x+3)/(x-5) is now 4(x-3)/(x+3)(x-5) correct?

Hero:

How do you figure that?

Kaihalojr:

because after we factor the first two fractions that's the only idea that seems fitting plus with fraction number 3 the numerator is x^2 - 25 which if squared is x - 5 so both (x - 5)s can cancel out.

Kaihalojr:

I say it's the only thing that fits because we must multiply

Hero:

You should write out your steps one by one here. Using the draw tool

Hero:

\(\dfrac{\dfrac{x + 3}{5x^2 -25x}}{\dfrac{28}{x^2 - 25}} = \) Write what the above equals after factoring the denominators